OpenStudy (anonymous):

In paragraph describe how to simplify each expression 16 1/2 sqrt 100

OpenStudy (anonymous):

16 1/2 = 16 + 1/2 = 32/2 +1/2 =33/2 or 16.5

OpenStudy (anonymous):

sqrt 100=lets 100 be x^2 then sqrt of 100 is = to sqrt x^2 which is equal to x, then x must be a multiple of it self rational or irrational. so therefore x must equal 10, or else it wouldnt be a square of itself because 10 is the only thing that works or else you would have X(Y) not X(X)